Choose an acid from the list below, that you could use to make a buffer having a pH of 8.91. After choosing an appropriate acid, propose concentrations of the two ingredients required for the buffer. Show that your proposed buffer will give a buffer pH of 8.91.

Choose an acid from the list below, that you couldPlease Show All Of Your Work!
Answer #1

Use the Henderson-Hasselbach equation:

pH = pKa + log ([A-]/[HA])

Assume that you're doing a simple buffer where [A-] = [HA], so log (1) = 0 and

pH = pKa = 8.91

pKa = -log(Ka)

Ka = 1.23 E-9. Now check the table for an approximated value.

The nearest is HBrO with Ka = 2.5 E-9, pKa = 8.6, so:

pH = pKa + log ([A-]/[HA])

log ([A-]/[HA]) = 8.91 - 8.6 = 0.31

[A-]/[HA] = 2.04

or: nA- = 2.04 nHA

For example: take 10 ml of 0.5 HBrO (0.005 moles). So you need 0.005*2.04 = 0.0102 moles of BrO-; for example, take 20.4 ml of 0.5 M BrO-

M HBrO = 0.005 / 0.0304 L = 0.164 M

M BrO- = 0.0102 / 0.0304 = 0.336

pH = 8.6 + log(0.336/0.164) = 8.91